please solve this question in details as the hidden part is “compounded monthly” .
For a Federal Credit Union that offers an interest rate of 8% per year, ly, deter- mine the nominal rate per 6 months.
Question : 4.9 For a Federal credit Union that offers as interest rate of 8% per year, compounded monthly, determine the nominal per 6 months .
Answer #1

Let, nominal rate per 6 month = R

Then,

(1+8%/12)^12 = (1+R)^2

Or

R = (1+8%/12)^6 - 1

R = 4.067% or 4.07%